OpenStudy (anonymous):

I know this isn't Physics this is Physical Science but I really need some help."A box 5 m long,and 2 m high applies 12 Pa of pressure to the table that is resting on.How much force does the box exert?"

OpenStudy (anonymous):

Pressure is defined as a force acting on an area. When something acts within a given region of space, this is normally expressed through division mathematically. So if pressure is a force acting within an area, then P = F / A. You're given P = 12 Pa You're only given the length and height of the box. I'm not sure if there's a width, so i'm not sure if the area will be correct. But you need to consider the area A that is in contact with the table. If that happens to be A = 5 * 2 = 10 m^2, then so be it So F = P * A = 12 * 10 = 120 N